OpenStudy (anonymous):

Don't Want Answer!!! Just want someone to help me, explain how to figure this out. How many solutions does the system of inequalities have? x-4y=12 and 5x-20y=60 a.one b.two c.infinetley many d. none

OpenStudy (loser66):

Question: they are equalities, right?

OpenStudy (anonymous):

oh no, I put inequalities!!! Sorry I meant to put equations! @Loser66

OpenStudy (loser66):

x-4y =12 5x-20y =60 right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

divide the second equation by 5 , what do you have?

OpenStudy (anonymous):

x-4y=12?

OpenStudy (loser66):

exactly the first one, right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

so, if (x1,y1) is the solution for the first one, it is the solution for the second one, too, right?

OpenStudy (anonymous):

yes, so the answer would be one solution?

OpenStudy (loser66):

not yet!!

OpenStudy (loser66):

if (x2,y2) is the solution for the first one, it is the solution for the second one, too if (x3,y3)----------------------------------------------------------- if (x4,y4)----------------------------------------------------------- and so on that means : any solution of the first one is the solution of the second one and to the first one, it has infinite solution ---> the answer is ?????

OpenStudy (anonymous):

infinetely many?

OpenStudy (loser66):

you got it by yourself, I don't answer it, right??? CONGRAT

OpenStudy (anonymous):

Oh my goodness. Thank you so much! I'm really bad at doing this and you just explained how to do it perfectly! c: